Goods Market Money Market
C = 250 + 0.8YD L = 0.25Y – 62.5i
YD = Y + TR – T Ms/P = 250
T = 100 + 0.25Y
I = 300 – 50i
G = 350; TR = 150

Goods market equilibrium condition: Y = C + I + G + X-M
Money market equilibrium condition: L = Ms/P
c) What are the equilibrium levels of income (Yo) and interest rate (io)
1
Expert's answer
2019-10-16T09:41:26-0400

c) The equilibrium levels of income (Yo) and interest rate (io) are:

Y = 250 + 0.8(Y + 150 - 100 - 0.25Y) + 300 - 50i + 350,

and

0.25Y – 62.5i = 250,

0.4Y = 940 - 50i and 0.25Y = 250 + 62.5i,

Y = 2350 - 125i and Y = 1000 + 250i,

2350 - 125i = 1000 + 250i,

375i = 1350,

io = 3.6,

Yo = 1000 + 250×3.6 = 1900.